# BayesFLo: Bayesian fault localization of complex software systems

### v2

@@ 5.2 Case Study 2: JMP Easy DOE Platform
1 Mode Guided, Flexible
2 N_Maximize_Responses 0, 1, 2, 3-5
3 N_Match_Target_Responses 0, 1, 2, 3-5
4 N_Minimize_Responses 0, 1, 2, 3-5
5 N_None_Responses 0, 1, 2, 3-5
6 N_Continuous_Factors 0, 1, 2, 3-5
7 N_Discrete_Numeric_Factors 0, 1, 2, 3-5
8 N_Categorical_Factors 0, 1, 2, 3-5
9 N_Discrete_Numeric_Levels 0, 1, 2, 3-5
10 N_Categorical_Levels 0, 1, 2, 3-5
11 Model_Type Main Effects, Main Effects Interact Corr, Main Effects Interact Uncorr, RSM
12 N_Extra_Runs 0, 2, 3, 4, 5-8
@@
Table 4: Input factors and their corresponding levels for the JMP Easy DOE case study.
Consider next the second case study on the fault localization of JMP’s Easy DOE Platform (JMP Statistical Discovery LLC,, 2023). This platform provides a popular end-to-end framework for experimental design in practical applications, from response and factor definition, model effects and run size specification, design generation, data analysis and results reporting. In testing this platform for deployment, however, the test engineers at JMP observed a failure in its design generation step: the generated design may have fewer runs than the specified run size. To pinpoint the root cause of this failure, the engineers investigated the $I=12$ factors in this module. Table 4 summarizes these factors and their corresponding levels. Here, the first factor controls the usage mode for Easy DOE: “Guided” guides the user through each step of the experimental design process, whereas “Flexible” offers a richer set of options for experienced users. The next four (factors 2 - 5) are response-related factors, which specify the number of responses and their associated goals. The following five (factors 6 - 10) then specify the types and quantities of experiment inputs. Finally, factor 11 indicates the type of statistical model, and factor 12 specifies the number of additional design points in addition to the default sample size from the software.
With this set-up, we construct a strength-2 covering array using the input factors and levels from Table 4, with a run size of $M=36$. Here, for each run, test engineers carefully examined the JMP output against the expected results based on a number of criteria to determine if such outputs align; if not, such a run is marked as a failure. Table 5 shows the corresponding test outcomes. Note that there is only one failure amongst the $M=36$ test runs; such a failure is the aforementioned mismatch in design run size. Using this data, Figure 5 shows the fault localization analysis from the JMP Covering Array module, which we again treat as the state-of-the-art. All eight combinations here are again tied in terms of their suspiciousness; the debugging of all such combinations can thus introduce high costs.
Factor
Run Number 1 2 3 4 5 6 7 8 9 10 11 12 Outcome
1 Guided 0 2 1 0 0 2 2 2 2 Main Effects 3 0
2 Guided 1 3-5 1 2 2 2 1 2 4-6 Main Effects Interact Corr 2 0
3 Guided 3-5 3-5 0 3-5 1 1 1 2 2 Main Effects Interact Corr 5-8 0
4 Flexible 1 0 3-5 3-5 2 2 1 4 4-6 Main Effects 3 0
5 Flexible 0 0 3-5 1 3-5 0 0 0 0 Main Effects Interact Uncorr 0 0
6 Flexible 2 0 3-5 2 1 2 3-5 4 2 RSM 2 0
7 Guided 1 1 0 0 2 1 0 4 0 RSM 0 0
8 Guided 2 0 2 0 3-5 3-5 1 3 4-6 Main Effects Interact Corr 5-8 0
9 Flexible 1 0 0 2 0 2 2 3 4-6 RSM 5-8 0
10 Guided 1 0 1 1 2 3-5 3-5 4 4-6 Main Effects 5-8 0
11 Guided 3-5 1 3-5 2 1 3-5 2 4 4-6 RSM 4 0
12 Flexible 0 3-5 1 3-5 0 3-5 0 4 0 Main Effects 4 0
13 Guided 3-5 2 2 3-5 1 1 0 3 0 Main Effects 3 0
14 Flexible 3-5 1 3-5 0 3-5 0 1 0 2 Main Effects Interact Uncorr 2 1
15 Flexible 0 2 0 3-5 2 2 1 2 4-6 RSM 2 0
16 Flexible 0 1 2 2 0 2 1 3 2 Main Effects 0 0
17 Guided 2 2 0 3-5 1 3-5 3-5 4 3 Main Effects Interact Corr 0 0
18 Guided 1 3-5 3-5 1 1 0 1 0 4-6 RSM 0 0
19 Flexible 1 1 2 1 1 3-5 2 4 2 RSM 0 0
20 Flexible 3-5 0 1 2 1 2 2 3 3 Main Effects 5-8 0
21 Flexible 3-5 1 1 3-5 0 0 3-5 0 4-6 RSM 5-8 0
22 Guided 2 0 3-5 2 1 1 0 2 0 Main Effects 2 0
23 Guided 3-5 2 0 1 2 2 1 3 2 Main Effects 4 0
24 Flexible 0 3-5 3-5 1 0 1 3-5 3 3 Main Effects 4 0
25 Flexible 2 3-5 0 2 1 0 2 0 4-6 Main Effects Interact Corr 3 0
26 Guided 2 1 1 1 3-5 3-5 3-5 2 3 RSM 3 0
27 Guided 0 0 2 0 1 0 1 0 3 Main Effects 4 0
28 Guided 0 1 2 0 0 1 3-5 4 4-6 Main Effects Interact Corr 2 0
29 Guided 1 3-5 1 3-5 3-5 1 2 2 3 Main Effects Interact Corr 4 0
30 Flexible 0 1 2 2 3-5 2 0 2 0 Main Effects 0 0
31 Guided 1 2 2 1 2 3-5 2 3 3 Main Effects 2 0
32 Flexible 1 2 0 2 3-5 0 2 0 2 Main Effects Interact Uncorr 4 0
33 Guided 0 2 3-5 1 3-5 2 0 4 0 Main Effects Interact Corr 5-8 0
34 Guided 2 3-5 2 0 2 0 2 0 0 Main Effects Interact Uncorr 4 0
35 Flexible 2 2 1 3-5 0 1 1 0 2 Main Effects Interact Uncorr 0 0
36 Guided 1 2 2 1 3-5 0 0 0 2 Main Effects Interact Corr 2 0
Table 5: The $M=36$-run CA design and test outcomes for the JMP Easy DOE case study. Here, an outcome of 0 indicates a passed test case, with 1 indicating a failed test case.
Next, we provide a careful prior specification for BayesFLo. Unlike the previous case study, here the test engineers have some prior domain knowledge that can be incorporated. In particular, from the nature of the considered failure (i.e., run size mismatch), the $I=12$ factors from Table 4 can be categorized into three categories of “suspiciousness”. The first category of low suspiciousness includes the four response-based factors (factors 2 - 5), which are least related to the design generation procedure. The second category of high suspiciousness includes the factor N_Extra_Runs, as this specifies the number of additional design runs and is thus most related to the considered failure. Finally, the last category of medium suspiciousness consists of the remaining seven factors. With this, we explore two choices of prior specifications:
  • Prior 1: To reflect these three categories, we set $p_{(i,j)}=0.02$ for factors with low prior suspicion, $p_{(i,j)}=0.04$ for factors with medium suspicion, and $p_{(i,j)}=0.16$ for the factor with high suspicion. This provides an informative prior that captures domain knowledge.
  • Prior 2: This adopts a less informative prior than Prior 1: it only integrates information on the response-based nature of factors 2 - 5. As such $p_{(i,j)}=0.02$ is set for such factors, and $p_{(i,j)}=0.04$ is set for remaining factors. This prior is included to investigate how additional domain knowledge can improve fault localization.
Top Suspicious Combinations (BayesFLo with Prior 1)
Factors Levels Posterior Probability
11, 12 Main Effects Interact Uncorr, 2 0.28
2, 12 3-5, 2 0.14
1, 5 Flexible, 0 0.03
2, 6 3-5, 3-5 0.03
2, 11 3-5, Main Effects Interact Uncorr 0.03
3, 11 1, Main Effects Interact Uncorr 0.03
4, 5 3-5, 0 0.02
2, 5 3-5, 0 0.02
Table 6: Top suspicious combinations from the BayesFLo analysis (with Prior 1) in the JMP Easy DOE case study, ranked by decreasing posterior root cause probabilities. Remaining combinations all have posterior probabilities less than $1.0\times 10^{-2}$. Here, the true root cause is bolded.
Top Suspicious Combinations (BayesFLo with Prior 2)
Factors Levels Posterior Probability
11, 12 Main Effects Interact Uncorr, 2 0.15
2, 12 3-5, 2 0.07
1, 5 Flexible, 0 0.07
2, 5 3-5, 3-5 0.07
2, 11 3-5, Main Effects Interact Uncorr 0.07
3, 11 1, Main Effects Interact Uncorr 0.07
4, 5 3-5, 0 0.04
2, 5 3-5, 0 0.04
Table 7: Top suspicious combinations from the BayesFLo analysis (with Prior 2) in the JMP Easy DOE case study, ranked by decreasing posterior root cause probabilities. Remaining combinations all have posterior probabilities less than $6.0\times 10^{-3}$. Here, the true root cause is bolded.
Figure 5: Top suspicious combinations from the JMP Covering Array analysis for the JMP Easy DOE case study, with its corresponding failure counts from test runs. Here, the true root cause is bolded.
Tables 6 and 7 show, for each prior choice, the top suspicious combinations from BayesFLo ranked by its posterior root cause probabilities. For both prior specifications, the combination with greatest posterior probability is the two-factor combination:
[…]
The probabilistic quantification of suspiciousness from BayesFLo also offers further insights. For the weakly-informative Prior 2, its top combination (which is the true root cause) has a posterior probability of 0.15, followed by seven combinations with posterior probabilities of 0.07 and 0.04; its remaining combinations all have probabilities less than $6.0\times 10^{-3}$. For the more informative Prior 1, its top combination (the true root cause) has a higher posterior probability of 0.28, followed by seven combinations with posterior probabilities of 0.14, 0.03 and 0.02; its remaining combinations all have probabilities less than $1.0\times 10^{-2}$. Thus, despite both priors identifying the true root cause as its most suspicious combination, the integration of more domain knowledge in Prior 1 provides heightened suspiciousness for this combination, which is as desired.
To contrast, the JMP analysis (Figure 5) is likewise more muddled for fault localization. Here, its top eight combinations are all tied, and the debugging of all such combinations to identify the true root cause can be costly. Its lack of probabilistic quantification on combination suspiciousness can similarly lead to a lack of confidence for fault localization. BayesFLo addresses such issues via a flexible probabilistic framework that integrates domain and structural knowledge: it can reduce testing costs by disentangling tied combinations to quickly pinpoint the true root cause for confident fault localization.
@@ 6 Conclusion
We proposed a new BayesFLo framework for Bayesian fault localization of complex software systems. Existing fault localization methods largely do not provide a probabilistic quantification of risk on potential root causes nor permit the integration of domain and/or structural knowledge from test engineers, which results many tied suspicious combinations and thus costly fault diagnosis. BayesFLo addresses such limitations via a new Bayesian framework on fault localization, which embeds the combination hierarchy and heredity principles (Lekivetz and Morgan,, 2021) to capture the structured nature of software root causes. One challenge is the computation of posterior root cause probabilities for BayesFLo, which can be infeasible even for small systems. We thus developed a new algorithmic framework for computing the desired posterior probabilities, leveraging recent tools from integer programming and graph representations. We then demonstrate the effectiveness of BayesFLo over the state-of-the-art in two case studies on the Traffic Alert Collision Avoidance System and the JMP Easy DOE platform.
Given promising results, there are many immediate avenues for future work. One direction is the use of the BayesFLo framework for sequential design of subsequent test sets. This adaptive testing of software, facilitated by the proposed Bayesian model, can greatly accelerate the discovery of bugs in complex systems. Another direction is the extension of BayesFLo for fault localization of systems with mixed (i.e., continuous and discrete) factors. Such a setting would be more complex, as it requires the probabilistic modeling of the fault response surface; recent work in Chen et al., (2022) appears to be useful for this goal. Finally, in the rare scenario where one encounters multiple failure types, a reasonable approach may be a Bayesian fault localization framework with categorical responses that can pinpoint root causes for different failure types; we will explore this in future work.
@@ Appendix A Proof for Proposition 1
$\begin{split}\mathcal{M}_{(\mathbf{i},\mathbf{j})}&=\{m=1,\cdots,M:y_{m}=1,(\mathbf{i},\mathbf{j})\in\mathcal{C}_{{\rm TF},m}\},\\ \mathcal{M}_{-(\mathbf{i},\mathbf{j})}&=\{m=1,\cdots,M:y_{m}=1,(\mathbf{i},\mathbf{j})\notin\mathcal{C}_{{\rm TF},m}\}.\\ \end{split}$
@@
$\begin{split}E_{\text{P}}&=\{\text{$Z_{c}=0$ for all $c\in\mathcal{C}_{\rm TP}$}\},\\ E_{(\mathbf{i},\mathbf{j})}&=\{\textup{for each $m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}$, there exists some $c\in\mathcal{C}_{{\rm TF},m}\setminus\mathcal{C}_{\rm TP}$ such that $Z_{c}=1$}\},\\ E_{-(\mathbf{i},\mathbf{j})}&=\{\textup{for each $m^{\prime}\in\mathcal{M}_{-(\mathbf{i},\mathbf{j})}$, there exists some $c\in\mathcal{C}_{{\rm TF},m^{\prime}}\setminus\mathcal{C}_{\rm TP}$ such that $Z_{c}=1$}\}.\end{split}$
@@ Appendix B Proof for Proposition 2
$E_{(\mathbf{i},\mathbf{j})}=\{\textup{for each $m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}$, there exists some $c\in\mathcal{C}_{{\rm TF},m}\setminus\mathcal{C}_{{\rm TP}}$ such that $Z_{c}=1$}\}.$
@@
$F_{(\mathbf{i},\mathbf{j})}=\{\{Z_{c}=1\textup{ for all }c\in\tilde{\mathcal{C}}\}\textup{, for at least one minimal cover $\tilde{\mathcal{C}}$ of $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$}\}.$
@@
Consider first $E_{(\mathbf{i},\mathbf{j})}\subseteq F_{(\mathbf{i},\mathbf{j})}$. This must be true, since if we collect all TF combinations $c$ with $Z_{c}=1$ from $E_{(\mathbf{i},\mathbf{j})}$, they contain at least one minimal cover $\tilde{\mathcal{C}}$, as all failures $m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ are covered. This suggests that every arbitrary element in $E_{(\mathbf{i},\mathbf{j})}$ is also an element in $F_{(\mathbf{i},\mathbf{j})}$. Consider next $F_{(\mathbf{i},\mathbf{j})}\subseteq E_{(\mathbf{i},\mathbf{j})}$. This must also be true, since if we take a minimal cover $\tilde{\mathcal{C}}$ of $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ with $Z_{c}=1$ for all $c\in\tilde{\mathcal{C}}$ from $F_{(\mathbf{i},\mathbf{j})}$, then for every $m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ there exists at least one $c\in\tilde{\mathcal{C}}$ that explains this failure. Thus every arbitrary element in $F_{(\mathbf{i},\mathbf{j})}$ is also an element in $E_{(\mathbf{i},\mathbf{j})}$. This proves the proposition.

### v3

@@ 5.2 Case Study 2: Vulnerable Road User Protection Tests
1 Daytime morning, day, evening, night
2 Haze/Fog no, yes
3 Street condition dry, wet, icy, snow, broken
4 Sky cloudy, no, clear
5 Rain no, yes
6 Reflection on road no, yes
7 Shadow on road no, yes
8 VRU type adult, child
9 VRU pose pedestrian, jogger, cyclist
10 VRU contrast to BG low, high
11 Speed slow, fast
12 Vehicle size small, medium, large
13 Lane number two, four
14 Camera no, yes
15 Direction same, opposite
@@
Table 5: Input factors and their corresponding levels for the VRU case study.
With this, we perform test runs using a strength-2 covering array with $M=23$ runs. These runs and their test outcomes are provided in Appendix C for brevity, where six runs resulted in failures. The BEN approach in Ghandehari et al. (2020) is then applied to the test outcomes, with its top-ranked suspicious combinations summarized in Table 6(a). Implementing the weighting approach in Lekivetz and Morgan (2018), however, comes with computational challenges. Recall from Section 2 that the computation of its normalized weights requires a complete enumeration of potential root cause scenarios, which can be infeasible when there are many failed test runs. Here, the six failed runs in our test set (see Appendix C) give rise to 54 2-factor potential root causes; a complete enumeration involves $2^{54}=1.80\times 10^{16}$ possible scenarios, which is infeasible to perform. We thus do not compare with the weighting approach benchmark in this case study.
As before, our BayesFLo approach is implemented under an informative and uninformative prior specification. For the informative prior, we adopt the baseline prior probability $p_{(i,j)}=1/38$, where its denominator $\sum_{i=1}^{I}J_{i}=38$ is the total number of levels over all $I$ factors. This baseline prior is assigned to the first ten factors (from the original system in Gladisch et al., 2020), and a higher prior $p_{(i,j)}=2/38$ is assigned to the last five factors, which are newly added and have not been thoroughly tested. For the uninformative prior, all 15 factors are assigned the baseline prior $p_{(i,j)}=1/38$.
Table 6 (b) and (c) show the suspicious 2-FCs from the informative and uninformative BayesFLo approaches, ranked by their posterior root cause probabilities. Here, all 3-FCs have lower probabilities than the lowest-ranked 2-FC, and are thus not reported for brevity. We see that the informative BayesFLo approach is able to pinpoint the true 2-factor root cause (17) as its most suspicious combination, with a posterior probability of 0.67. The next four 2-FCs have posterior probabilities between 0.28 and 0.44, and the remaining 2-FCs have probabilities less than 0.25. As before, BayesFLo is preferable to the BEN approach (see Table 6(a)) from a debugging perspective: it finds the true root cause as its top combination, whereas BEN requires the investigation of four additional combinations before finding the root cause. There are two reasons for this: (i) BEN does not incorporate prior domain knowledge in its analysis, and (ii) BEN relies on single-factor components for evaluating its suspiciousness metrics (see Section 2), which may obscure the identification of multi-factor root causes. Comparing the informative BayesFLo with its uninformative counterpart, we again see that the incorporation of prior domain knowledge can improve fault localization: such knowledge helps break the tie of the two top 2-FCs in the uninformative analysis, allowing BayesFLo to pinpoint the true root cause as its top combination.
Finally, this probabilistic BayesFLo analysis also helps guide actionable decisions for confident fault localization. From the BEN analysis (see Table 6(a)), it is again difficult to plan the subsequent debugging effort, since it is not clear which combinations can be confidently ignored. BayesFLo can give a clearer picture. With a significance level of $\alpha=0.05$, the informative BayesFLo analysis suggests that 39 combinations need to be further investigated; the last 15 2-FCs and all 3-FCs (with probabilities less than $\alpha$) can be confidently ignored. Note that the set of combinations to debug here is larger than that from the earlier TCAS case study. This is quite intuitive: there are considerably more failed test runs here, which should result in heightened suspicion and a larger debugging effort. Our probabilistic BayesFLo analysis thus captures this intuition to guide an effective planning of debugging efforts.
Type Count Rank
2-FC 4 1-4
2-FC 1 5
2-FC 49 6-54
(a) BEN Phase 1.
Type Count Probability
2-FC 1 0.50
2-FC 1 0.50
2-FC 2 0.33
2-FC 50 0.05-0.25
(b) BayesFLo (uninformative).
Type Count Probability
2-FC 1 0.67
2-FC 1 0.44
2-FC 1 0.33
2-FC 1 0.32
2-FC 2 0.28
2-FC 33 0.05-0.25
2-FC 15 <<0.05
(c) BayesFLo (informative).
Table 6: Suspicious 2-FCs ranked using the BEN approach (Phase 1) and BayesFLo (with uninformative and informative priors) for the VRU case study. Combinations at the top are deemed more suspicious and are investigated first. The true root cause is bolded.
@@ 6 Conclusion
We proposed a new BayesFLo framework for Bayesian fault localization of complex software systems. Compared to existing methods, BayesFLo can accelerate the identification of root causes with a small number of test runs, and provide a principled quantification of probabilistic risk that can guide subsequent debugging efforts. This is achieved via a new Bayesian fault localization model, which embeds the combination hierarchy and heredity principles (Lekivetz and Morgan, 2021) to capture the structured nature of software root causes. One challenge is the computation of posterior root cause probabilities for BayesFLo, which can be infeasible even for small systems. We thus developed a new algorithmic framework for computing the desired posterior probabilities, leveraging recent tools from integer programming and graph representations. We then demonstrate the effectiveness of BayesFLo over the state-of-the-art in case studies on the Traffic Alert Collision Avoidance System and the Vulnerable Road User protection tests.
Building on these promising results, there are many immediate avenues for future work. One direction is the use of the BayesFLo framework for sequential design of subsequent test sets. This adaptive testing of software, facilitated by the proposed Bayesian model, can accelerate the discovery of bugs in complex systems; recent work in Chen et al. (2022) appears to be useful for this goal. Another direction is the extension of BayesFLo for fault localization of systems with mixed (i.e., continuous and discrete) factors. Such a setting would be more complex, as it requires the probabilistic modeling of the fault response surface. Finally, in the rare scenario where one encounters multiple failure types, a reasonable approach may be a Bayesian fault localization framework with categorical responses that can pinpoint root causes for different failure types; we will explore this in future work.
@@ Appendix A Proof for Proposition 1
[…]
@@
[…]
@@ Appendix B Proof for Proposition 2
[…]
@@
[…]
@@
Consider first $E_{(\mathbf{i},\mathbf{j})}\subseteq F_{(\mathbf{i},\mathbf{j})}$. This must be true, since if we collect all TF combinations $c$ with $Z_{c}=1$ from $E_{(\mathbf{i},\mathbf{j})}$, they contain at least one minimal cover $\tilde{\mathcal{C}}$, as all failures $m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ are covered. This suggests that every arbitrary element in $E_{(\mathbf{i},\mathbf{j})}$ is also an element in $F_{(\mathbf{i},\mathbf{j})}$. Consider next $F_{(\mathbf{i},\mathbf{j})}\subseteq E_{(\mathbf{i},\mathbf{j})}$. This must also be true, since if we take a minimal cover $\tilde{\mathcal{C}}$ of $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ with $Z_{c}=1$ for all $c\in\tilde{\mathcal{C}}$ from $F_{(\mathbf{i},\mathbf{j})}$, then for every $m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ there exists at least one $c\in\tilde{\mathcal{C}}$ that explains this failure. Thus every arbitrary element in $F_{(\mathbf{i},\mathbf{j})}$ is also an element in $E_{(\mathbf{i},\mathbf{j})}$. This proves the proposition. ∎
Appendix C Test design and outcomes for the VRU case study
Factor
Run Number 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 Outcome
1 4 1 1 3 2 2 1 1 3 1 1 3 2 2 2 0
2 4 1 2 2 2 2 2 1 2 2 2 2 2 1 2 1
3 3 2 3 3 1 2 1 1 2 2 1 3 1 1 2 0
4 1 1 1 1 1 2 1 2 1 2 1 1 1 1 2 0
5 2 2 5 3 1 1 2 1 2 2 2 2 2 1 2 0
6 4 2 4 3 1 2 1 2 3 2 1 1 2 2 1 0
7 3 2 5 2 2 2 1 1 3 1 2 3 2 1 1 1
8 2 2 1 2 1 2 2 2 2 1 1 2 1 1 1 0
9 4 1 3 3 1 1 2 2 1 1 2 2 1 2 1 0
10 1 1 5 3 2 1 1 2 2 1 1 1 2 2 1 0
11 3 2 1 1 1 1 2 1 1 1 2 2 2 2 1 0
12 4 2 5 1 2 1 2 2 1 2 1 3 1 1 1 0
13 2 1 3 1 1 1 2 1 2 2 2 1 1 2 1 0
14 1 2 4 2 2 2 1 1 3 1 2 2 2 1 2 1
15 2 1 4 2 2 2 2 2 1 1 2 3 1 1 2 1
16 2 2 2 1 1 2 1 2 3 1 1 1 1 2 1 0
17 1 1 2 2 2 2 2 2 1 2 2 3 2 2 2 1
18 1 1 3 2 1 1 2 2 3 1 2 1 2 2 2 0
19 3 1 4 3 1 1 1 2 2 1 2 1 1 2 2 0
20 3 1 2 2 2 2 1 2 3 1 2 1 2 2 1 1
21 2 2 2 3 1 1 1 2 3 1 1 1 1 2 1 0
22 4 2 4 1 1 2 1 2 3 2 1 1 2 2 1 0
23 2 1 3 1 2 1 2 1 2 2 2 1 1 2 1 0
Table 7: The $M=23$-run CA design and test outcomes for the VRU case study. Here, an outcome of 0 indicates a passed test case, with 1 indicating a failed test case.